stephanie sold 9 apple pies and 10 pumpkin for 277 dollars. Krystal sold 2 apple and 5 pumpkin for 106 dollars. whats the cost of one apple pie and one pumpkin pie?
Answer by FrankM(1040)

9A+10P=277
2A+5P=106
2 x the second equation
4A+10P=212
subtract from first
5A=65
A=13
2(13)+5P=106
26+5P=106
5P=80
P=16
add A+P